OpenStudy (rikkie):

The recursive formula for an arithmetic sequence is given below. What is the third term in the sequence? a1=-9 an=3+a(n-1) a. -6 b. -3 c. 3 d. 0

ganeshie8 (ganeshie8):

\(\large \mathbb{a_1 = -9}\) \(\large \mathbb{a_n = 3 + a_{n-1}}\)

ganeshie8 (ganeshie8):

first find \(a_2\) put \(n=2\) in the given recursive formula

ganeshie8 (ganeshie8):

\\(\large \mathbb{a_2 = 3 + a_{2-1}}\) \\(\large \mathbb{~~~ = 3 + a_{1}}\)

ganeshie8 (ganeshie8):

plugin \(a_1\) value above, similarly find the third term, \(a_3\), using \(a_2\) value